OpenStudy (anonymous):

If (u,v) = u + v/u + v, find F(1/u, 1/v) + F(u, v).

OpenStudy (jdoe0001):

$$ \text{if } \ \ f(u,v) = \cfrac{u+v}{u+v}\\ \text{find}\\ f\pmatrix{\frac{1}{u}, \frac{1}{v}}+f(u,v) \huge ? $$

OpenStudy (anonymous):

Yes.

OpenStudy (jdoe0001):

hmmm, are just expecting the expand, because is quite unnecessary :/

OpenStudy (jdoe0001):

a/a = 1 so same/same =1 and \(\cfrac{u+v}{u+v} = 1\)

OpenStudy (jdoe0001):

the 2nd function with the reciprocals, simple make the variables a fraction, but you'd still remain with \(\large \frac{same}{same} \)

OpenStudy (jdoe0001):

$$ \text{if } \ \ f(u,v) = \cfrac{u+v}{u+v} = 1\\ \text{then}\\ f\pmatrix{\frac{1}{u},\frac{1}{v}} = \cfrac{ \frac{1}{u}+\frac{1}{v} }{ \frac{1}{u}+\frac{1}{v} } = 1 $$

OpenStudy (anonymous):

Oops, there has been a mistake. f(u, v) = u+v/u-v sorry :)

OpenStudy (anonymous):

using slope?

OpenStudy (anonymous):

\[f \left( u,v \right)=\frac{ u+v }{ u-v }\] \[f \left( \frac{ 1 }{ u },\frac{ 1 }{ v } \right)=\frac{ \frac{ 1 }{u }+\frac{ 1 }{ v } }{\frac{ 1 }{ u }-\frac{ 1 }{ v } }\] \[=\frac{ \frac{ v+u }{uv } }{ \frac{ v-u }{ uv } }\] \[=\frac{ v+u }{uv}*\frac{ uv }{ v-u }\] \[=-\frac{ u+v }{u-v }=-f \left( u,v \right)\] \[Hence f(u,v)+f \left( \frac{ 1 }{u },\frac{ 1 }{ v } \right)=0\]

OpenStudy (anonymous):

The exact question is: If (u, v)= u+v/u-v, find F(1/u, 1/v) + F(u, v).

OpenStudy (jdoe0001):

see surjithayer's line above :)

OpenStudy (whpalmer4):

@raechelvictoria look at all the effort that could have been saved if you typed the problem accurately in the first place, or looked carefully when jdoe0001 asked you if the typeset version was correct?

OpenStudy (jhannybean):

Yep, he's got the right idea.

OpenStudy (whpalmer4):

Be as careful as you want us to be in helping you :-)

OpenStudy (anonymous):

@whpalmer4 @jdoe0001 I'm really really sorry :(((((((( thank you for your efforts in answering my question :)

OpenStudy (jdoe0001):

yw